# Tensor products of bimodules and bimodule pairs over monoids

### v2

@@ 2.3.1. Homomorphisms of pairs
We consider $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas $\mathcal{A}$ and $(\mathcal{T}_{1}^{\prime},\mathcal{T}_{2}^{\prime})$-bimagmas $\mathcal{A}^{\prime}$, and functions $f:\mathcal{A}\to\mathcal{A}^{\prime}$ together with module homomorphisms $f_{\mathcal{T}_{i}}:\mathcal{T}_{i}\to\mathcal{T}_{i}^{\prime}.$ (When $\mathcal{A}$ is left weakly admissible then $f_{\mathcal{T}_{1}}$ is the restriction of $f$ to $\mathcal{T}_{1}.$ Likewise on the right.)
@@
    The support $\operatorname{supp}(f)$ of a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ is $\{b\in\mathcal{A}:f(b)\neq\iota^{\prime}\}.$
@@
    A bimagma homomorphism (also just called homomorphism) $f:\mathcal{A}\to\mathcal{A}^{\prime}$ is a multiplicative map satisfying $f(b_{1}*b_{2})=f(b_{1})*f(b_{2}),$ $\forall b_{i}\in\mathcal{A}.$
  4. (iv)
@@
For $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas, $\operatorname{Hom}(\mathcal{A},\mathcal{A}^{\prime})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma under the operation $(f_{1}*f_{2})(b)=f_{1}(b)*f_{2}(b)$ and the actions $af:b\mapsto af(b)$ and $fa:b\mapsto f(b)a$.
@@
Suppose $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ are $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepairs.
@@
    Given a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ and any homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ , $(\mathcal{A}^{\prime},f(\mathcal{A}_{0}))$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair, and $f$ can be viewed as a paired homomorphism $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},f(\mathcal{A}_{0}))$.
@@
    For any $(\mathcal{T}_{1},\mathcal{T}_{2})$-biomodule homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ where $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair, there is a prepair $(\mathcal{A},\mathcal{A}_{0})$ where $\mathcal{A}_{0}=\{b\in\mathcal{A}:f(b)\in\mathcal{A}_{0}^{\prime}\},$ and then $f$ is a paired homomorphism.
@@
2.4. Pre-orders and surpassing relations on pairs
@@
    A surpassing relation on a $\mathcal{T}$-magma $\mathcal{A}$ is a pre-order satisfying the following:
@@
In a prepair $(\mathcal{A},\mathcal{A}_{0})$ with a pre-order, $b\preceq b*c$ and $b\preceq c*b$ for all $b\in\mathcal{A}$ and $c\in\mathcal{A}_{0}$.
@@
Lemma 2.19 ([11, Lemma 2.11]).
If $b_{1}\preceq b_{2}$ holds in a pre-ordered prepair with a negation map, then $b_{2}(-)b_{1}\succeq\iota$ and $b_{1}(-)b_{2}\succeq\iota.$
@@
2.4.1. $\preceq$-morphisms
We insert the surpassing relation into our categories.
@@ 2.4.1. $\preceq$-morphisms
Let $(\mathcal{A},\mathcal{A}_{0})$, $(\mathcal{A}^{\prime},\mathcal{A}_{0}^{\prime})$ be prepairs.
@@
    When $\mathcal{A}$ also has a pre-order $\preceq$, a $\preceq$-morphism, (analogous to “colax morphism” in [20]) is a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying the following conditions:
@@
Given an 1:1 multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ where $\mathcal{A}$ is a $\mathcal{T}$-magma and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a pair with surpassing relation $\preceq$, the pair $(\mathcal{A},\mathcal{A}_{0})$ of Lemma 2.16 has a surpassing relation given by $b_{1}\preceq b_{2}$ when $f(b_{1})\preceq f(b_{2}),$ and then $f$ becomes a $\preceq$-morphism.
@@
2.5. Negation maps
At times we can define the negation map, the mainstay of [24]. A negation map ()(-)( - ) on a $\mathcal{T}$-module $(\mathcal{A},\mathcal{A}_{0})$ is an additive automorphism ()(-)( - ) of order $\leq 2$ also defined on $\mathcal{T}$, such that
@@ 2.5. Negation maps
We write $b_{1}(-)b_{2}$ for $b_{1}+((-)b_{2}).$ Thus $b^{\circ}=b(-)b,$ and $\mathcal{A}_{0}$ contains the set $\mathcal{A}^{\circ}=\{b^{\circ}:b\in\mathcal{A}\}.$ Often $\mathcal{A}_{0}=\mathcal{A}^{\circ}.$
@@
Hence, perhaps surprisingly, if $b_{1}\preceq b_{2},$ then $(-)b_{1}\preceq(-)b_{2}$.
@@ 2.6. Bimodules
We pinpoint the additively associative case.
@@
    A left $\mathcal{T}$-module $\mathcal{M}$ is a left $\mathcal{T}_{\iota}$-magma where * is also abelian and associative, in which case we write $+$ in place of *, and $\mathbb{0}$ instead of $\iota$.
@@
    A $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimodule is a $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimagma $(\mathcal{M},+)$ which is also abelian and associative.
@@
Proof.
Standard. ∎
Just as in classical ring theory and module theory, the category theory for semialgebras differs from the category theory for bimodules, we want to differentiate between $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimodules (resp. bimagmas) and $\mathcal{T}$-semialgebras (resp. pre-semialgebras), which we now define.
@@
    A $\mathcal{T}$-pre-semialgebra is a $(\mathcal{T},\mathcal{T})$-bimagma (with respect to *) where $\mathcal{T}$ is $\mathcal{A}$-central, which also has a second operation which we denote as $\cdot$, with $a_{1}(b_{1}\cdot b_{2})=(a_{1}b_{1})\cdot b_{2}$ and $(b_{1}\cdot b_{2})a_{2}=b_{1}\cdot(b_{2}a_{2})$ for $a_{i}\in\mathcal{T},$ $b_{1},b_{2}\in\mathcal{A}.$ In particular $a\cdot b=ab$ and $b\cdot a=ba$ for all $a\in\mathcal{T}$ and $b\in\mathcal{A}.$
@@
    A $\mathcal{T}$-semialgebra is a $\mathcal{T}$-pre-semialgebra which is a semiring, where * is the usual addition.
@@
    A pre-semialgebra $\mathcal{A}$ with a surpassing relation $\preceq$ is $\preceq$-distributive if $b(b_{1}*b_{2})\preceq bb_{1}\,*\,b_{2}b_{2}$ and $(b_{1}*b_{2})b\preceq b_{1}b\,*\,b_{2}b$ for all $b,b_{i}\in\mathcal{A}.$
@@ 2.6.1. Bimodule-theoretic notions for prepairs
The following notions are difficult to notate for bimagmas (which need not be associative), so we introduce them only for bimodules.
@@ 2.6.2. Weak morphisms
For prepairs $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0}),$ define $\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})_{0}$ to be the weak morphisms for which $f(\mathcal{A})\subseteq\mathcal{A}^{\prime}_{0}.$
@@
The main verification for (ii) is that the sum $f_{1}+f_{2}$ of weak morphisms is a weak morphism. If $\sum b_{i}\in\mathcal{A}_{0},$ then $(f_{1}+f_{2})(\sum b_{i})=f_{1}(\sum b_{i})+f_{2}(\sum b_{i})\in\mathcal{A}_{0}+\mathcal{A}_{0}\subseteq\mathcal{A}_{0}.$
@@
Every $\preceq$-morphism of prepairs is a weak morphism.
@@ 3. Hypermagmas and hyperpairs
    A hypermagma $(\mathcal{H},*)$ is a set with a binary operation :×$*:\mathcal{H}\times\mathcal{H}\to\mathcal{P}(\mathcal{H})$, permitting the product of two elements to be the empty set [^4] Customarily one requires $a_{1}*a_{2}\in{\mathcal{P}}^{\sharp}(\mathcal{H})$ for all $a_{1},a_{2}\in\mathcal{H}.$ But one could formally adjoin $\emptyset$, so the Nakamura-Reyes definition subsumes the customary definition, and we shall need it in what follows. $\emptyset$ could be interpreted as “undefined.”, together with a left and right $\mathcal{H}$-action whose natural extension to $\mathcal{P}(\mathcal{H})$ makes $(\mathcal{P}(\mathcal{H}),*)$ an $\mathcal{H}$-bimagma when we define, for $S_{1},S_{2}\in\mathcal{P}(\mathcal{H}),$
@@
    From now on we follow the customary use of $\boxplus$ and $\mathbb{0}$ instead of * and $\iota$, when the operation is associative. We call $\boxplus$ “hyperaddition.”
@@
    A hypernegative of an element $a$ in $(\mathcal{H},\boxplus,\mathbb{0})$ (if it exists) is an element “$-a$” for which $\mathbb{0}\in a\boxplus(-a)$ and $\mathbb{0}\in(-a)\boxplus a$.
@@
    If the hypernegative $-\mathbb{1}$ exists in $\mathcal{H},$ then we define $e=\mathbb{1}\boxplus(-\mathbb{1}).$
@@
    A hypergroup is a hypersemigroup $(\mathcal{H},\boxplus,\mathbb{0})$ for which every element $a$ has a unique hypernegative denoted $-a$, whereby, for all $a_{i}\in\mathcal{H},$
@@
    A hypersemiring (resp. hyperring) is an abelian [^6] In [20] this is called “canonical.” hypersemigroup (resp. hypergroup) $(\mathcal{H},\boxplus,\mathbb{0})$, providing $\mathcal{P}(\mathcal{H})$ with a natural elementwise multiplication, distributing over hyperaddition, making $\mathcal{P}(\mathcal{H})$ a pre-semialgebra pair over $\mathcal{T}=\mathcal{H}\setminus\{\mathbb{0}\}$ under the action $aS=\{as:s\in S\}$. [^7] In general $\mathcal{P}(\mathcal{H})$ is not distributive, but does satisfy $(\boxplus_{i}S_{i})(\boxplus_{j}S_{j}^{\prime})\subseteq(\boxplus_{i}S_{i}S_{j}^{\prime})$, cf. [18, Proposition 1.1].
@@
    A hypergroup $\mathcal{H}$ is a hyperfield if $\mathcal{H}\setminus\{\mathbb{0}\}$ is a multiplicative group, with $\mathbb{0}a=\mathbb{0}$ for all $a\in\mathcal{H}.$
@@
    If it exists, the hyperzero of a hypersemigroup is unique, by the familiar elementary argument. We write $\mathcal{H}$ for the $(\mathcal{H},\boxplus,\mathbb{0})$ in the rest of this remark.
@@
    $\mathcal{H}$ gives rise to a $\mathcal{H}$-magma $\mathcal{P}(\mathcal{H}),$ with hyperaddition given by
@@
    Thus $\emptyset$ is the $\infty$ element in the sense of Definition 2.1.
@@
    $\mathcal{H}$ is a weakly admissible $\mathcal{T}$-bimagma in case
@@
Say a subset $S\subseteq\mathcal{P}(\mathcal{H})$ is weakly neutral if $a\in a+S$ for each $a\in\mathcal{H}$. The weakly neutral subsets comprise a submagma of $\mathcal{P}(\mathcal{H})$.
@@
Definition 3.4.
In Remark 3.2(v), the sub-pair of the pair $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$ generated by $\mathcal{H}$ is called the hyperpair of $\mathcal{H}.$ [^8] This is the definition in [2], but although it could lead to difficulties if $\mathcal{P}(\mathcal{H})$ is not distributive, because then the $\mathcal{H}$-submagma of the power set $\mathcal{P}(\mathcal{H})$ spanned by $\mathcal{H}$ need not be closed under multiplication. Alternatively, one could take $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$, but it need not be admissible.
Remark 3.5.
  1. (i)
    Any hyperpair has the important surpassing relation $\subseteq$,⊆ , i.e., $S_{1}\preceq S_{2}$ when $S_{1}\subseteq S_{2}.$ Of course $\mathcal{P}(\mathcal{H})$ is a lower semilattice with respect to ()($\subseteq$)( ⊆ ), but not ${\mathcal{P}}^{\sharp}(\mathcal{H}).$
  2. (ii)
    More generally, if $\mathcal{H}$ has a surpassing relation $\preceq$, then $\mathcal{P}(\mathcal{H})$ has the surpassing relation $S_{1}\preceq S_{2}$ if for each $s_{1}\in S_{1}$ there is $s_{2}\in S_{2}$ for which $s_{1}\preceq s_{2}.$
  3. (iii)
    For hypermagmas $\mathcal{H}$, the condition that $b_{1}\subseteq b_{2}$ implies $f(b_{1})\subseteq f(b_{2})$ is automatic on $\mathcal{P}(\mathcal{H})$.
  4. (iv)

### v4

@@ 2.2.2. Homomorphisms of pairs
We consider $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas $\mathcal{A}$ and $(\mathcal{T}_{1}^{\prime},\mathcal{T}_{2}^{\prime})$-bimagmas $\mathcal{A}^{\prime}$, and functions $f:\mathcal{A}\to\mathcal{A}^{\prime}$ together with monoid homomorphisms $f_{\mathcal{T}_{i}}:\mathcal{T}_{i}\to\mathcal{T}_{i}^{\prime}.$
@@
    A bimagma homomorphism (also just called a homomorphism) $f:\mathcal{A}\to\mathcal{A}^{\prime}$ is a multiplicative map satisfying $f(b_{1}*b_{2})=f(b_{1})*f(b_{2}),$ $\forall b_{i}\in\mathcal{A}.$
@@
For $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas $\mathcal{A}$ and $\mathcal{A}^{\prime}$, $\operatorname{Hom}(\mathcal{A},\mathcal{A}^{\prime})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma under the operation $(f_{1}*f_{2})(b)=f_{1}(b)*f_{2}(b)$ and the actions $af:b\mapsto af(b)$ and $fa:b\mapsto f(b)a$.
@@
Suppose $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ are prepairs.
@@
    Given a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ and any homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ , $(\mathcal{A}^{\prime},f(\mathcal{A}_{0}))$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair, and then $f$ is a paired homomorphism $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},f(\mathcal{A}_{0}))$.
@@
    For any $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ where $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair, there is a prepair $(\mathcal{A},\mathcal{A}_{0})$ where $\mathcal{A}_{0}=\{b\in\mathcal{A}:f(b)\in\mathcal{A}_{0}^{\prime}\},$ and then $f$ is a paired homomorphism.
@@
2.2.3. Pre-orders and surpassing relations
@@
    A surpassing relation on a weakly admissible $\mathcal{T}$-bimagma $\mathcal{A}$ is a pre-order satisfying the following:
@@
In a prepair $(\mathcal{A},\mathcal{A}_{0})$ with a surpassing relation, $b\preceq b*c$ and $b\preceq c*b$ for all $b\in\mathcal{A}$ and $c\in\mathcal{A}_{0}$.
@@
We insert the preorder into our categories.
@@
Let $\mathcal{A},\mathcal{A}^{\prime}$ be pre-ordered $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas.
@@
    A $\preceq$-morphism, (analogous to “colax morphism” in [20]) is a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying the following conditions:
@@
Given an 1:1 multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ where $\mathcal{A}$ is a $\mathcal{T}$-bimagma and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a pair with surpassing relation $\preceq$, the pair $(\mathcal{A},\mathcal{A}_{0})$ of Lemma 2.15(ii) has a surpassing relation given by $b_{1}\preceq b_{2}$ when $f(b_{1})\preceq f(b_{2}),$ and then $f$ becomes a $\preceq$-morphism.
@@
2.2.4. Negation maps
At times we can define the negation map, the mainstay of [24]. A negation map ()(-) on a $\mathcal{T}$-module $(\mathcal{A},\mathcal{A}_{0})$ is an additive automorphism ()(-) of order $\leq 2$ also defined on $\mathcal{T}$, such that
@@ 2.2.4. Negation maps
We write $b_{1}(-)b_{2}$ for $b_{1}*((-)b_{2}).$ Thus $b^{\circ}=b(-)b,$ and $\mathcal{A}_{0}$ contains the set $\mathcal{A}^{\circ}=\{b^{\circ}:b\in\mathcal{A}\}.$ Often $\mathcal{A}_{0}=\mathcal{A}^{\circ}.$
@@
Hence, perhaps surprisingly, if $b_{1}\preceq b_{2},$ then $(-)b_{1}\preceq(-)b_{2}$. Hence, if $b_{1}\preceq b_{2}$ then $b_{2}(-)b_{1}\succeq\iota$ and $b_{1}(-)b_{2}\succeq\iota,$ cf. [11, Lemma 2.11].
@@ 2.3. Bimodules
We pinpoint the case of greatest interest.
@@
    A left $\mathcal{T}$-module $\mathcal{M}$ is a left $\mathcal{T}_{\iota}$-magma where * is also commutative and associative, in which case we always write ++ in place of *, and $\mathbb{0}$ instead of $\iota$.
@@
    A $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimodule is a $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimagma which is a left $\mathcal{T}_{1}$-module and a right $\mathcal{T}_{2}$-module.
@@
Just as the category theory for (semi)algebras differs from the category theory for bimodules, we want to differentiate between $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimodules (resp. bimagmas) and $\mathcal{T}$-semialgebras (resp. pre-semialgebras), which we now define.
@@
    A $\mathcal{T}$-pre-semialgebra is a $\mathcal{T}$-bimodule $(\mathcal{A},+)$, which also has a second operation which we denote as $\cdot$, with $a_{1}(b_{1}\cdot b_{2})=(a_{1}b_{1})\cdot b_{2}$ and $(b_{1}\cdot b_{2})a_{2}=b_{1}\cdot(b_{2}a_{2})$ for $a_{i}\in\mathcal{T},$ $b_{1},b_{2}\in\mathcal{A}.$ In particular, when $\mathcal{A}$ is weakly admissible, $a\cdot b=ab$ and $b\cdot a=ba$ for all $a\in\mathcal{T}$ and $b\in\mathcal{A}.$
@@
    A $\mathcal{T}$-semialgebra is a $\mathcal{T}$-pre-semialgebra which is a semiring.
@@
    A pre-semialgebra $\mathcal{A}$ with a surpassing relation $\preceq$ is $\preceq$-distributive if $b(b_{1}+b_{2})\preceq bb_{1}\,+\,b_{2}b_{2}$ and $(b_{1}+b_{2})b\preceq b_{1}b\,+\,b_{2}b$ for all $b,b_{i}\in\mathcal{A}.$
@@ 2.3.1. Bimodule-theoretic notions for prepairs
The following notions are difficult to notate for bimagmas (which need not be associative under *), so we introduce them only for bimodules.
@@ 2.3.2. Weak morphisms
For bimodule prepairs $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0}),$ define $\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})_{0}$ to be the weak morphisms for which $f(\mathcal{A})\subseteq\mathcal{A}^{\prime}_{0}.$
@@
(i) is as in Lemma 2.13.
The main verification for (ii) is that the sum $f_{1}+f_{2}$ of two weak morphisms is a weak morphism. If $\sum b_{i}\in\mathcal{A}_{0},$ then $(f_{1}+f_{2})(\sum b_{i})=f_{1}(\sum b_{i})+f_{2}(\sum b_{i})\in\mathcal{A}_{0}+\mathcal{A}_{0}\subseteq\mathcal{A}_{0}.$
@@
Every $\preceq$-morphism of bimodule prepairs is a weak morphism.
@@ 3. Hypermagmas and hyperpairs
    A hypermagma $(\mathcal{H},*)$ is a set with a binary operation $\mathcal{H}\times\mathcal{H}\to\mathcal{P}(\mathcal{H})$, permitting the product of two elements to be the empty set [^4] Customarily one requires $a_{1}*a_{2}\in{\mathcal{P}}^{\sharp}(\mathcal{H})$ for all $a_{1},a_{2}\in\mathcal{H}.$ But one could formally adjoin $\emptyset$, so the Nakamura-Reyes definition subsumes the customary definition, and we shall need it in what follows. $\emptyset$ could be interpreted as “undefined.”, together with a left and right $\mathcal{H}$-action whose natural extension to $\mathcal{P}(\mathcal{H})$ makes $(\mathcal{P}(\mathcal{H}),*)$ an $\mathcal{H}$-bimagma when we define, for $S_{1},S_{2}\in\mathcal{P}(\mathcal{H}),$
@@
    From now on we follow the customary use of $\boxplus$ and $\mathbb{0}$ instead of * and $\iota$, to denote that the operation is associative. We call $\boxplus$ “hyperaddition.”
@@
    A hypernegative of an element $a$ in $(\mathcal{H},\boxplus,\mathbb{0})$ (if it exists) is an element “a-a” for which $\mathbb{0}\in a\boxplus(-a)$ and $\mathbb{0}\in(-a)\boxplus a$.
@@
    If the hypernegative $\mathbb{1}$ exists in $\mathcal{H},$ then we define $e=\mathbb{1}\boxplus(-\mathbb{1}).$
@@
    A hypergroup is a hypersemigroup $(\mathcal{H},\boxplus,\mathbb{0})$ for which every element $a$ has a unique hypernegative denoted a-a, whereby, for all $a_{i}\in\mathcal{H},$
@@
    A hypersemiring (resp. hyperring) is an abelian [^6] In [20] this is called “canonical.” hypersemigroup (resp. hypergroup) $(\mathcal{H},\boxplus,\mathbb{0})$, providing $\mathcal{P}(\mathcal{H})$ with a natural elementwise multiplication, distributing over hyperaddition, making $\mathcal{P}(\mathcal{H})$ a pre-semialgebra pair over $\mathcal{T}_{\iota}=\mathcal{H}$ under the action $aS=\{as:s\in S\}$. [^7] Thus $\mathbb{0}a=\mathbb{0}$ for all $a\in\mathcal{H}.$ Although not necessarily distributive, $\mathcal{P}(\mathcal{H})$ but satisfies $(\boxplus_{i}S_{i})(\boxplus_{j}S_{j}^{\prime})\subseteq(\boxplus_{i}S_{i}S_{j}^{\prime})$, cf. [18, Proposition 1.1].
@@
    A hypergroup $\mathcal{H}$ is a hyperfield if $\mathcal{H}\setminus\{\mathbb{0}\}$ is a multiplicative group.
@@
Although [20] uses hypermagmas, we work mostly with hypersemigroups and write $\mathcal{H}$ for $(\mathcal{H},\boxplus,\mathbb{0})$.
@@
    If it exists, the hyperzero of a hypersemigroup is unique, by the familiar elementary argument.
@@
    $\mathcal{H}$ gives rise to a $\mathcal{H}$-module $\mathcal{P}(\mathcal{H}),$ with hyperaddition given by
@@
    $\mathcal{P}(\mathcal{H})$ is a weakly admissible $\mathcal{T}$-bimodule in case
@@
Definition 3.3.
In Remark 3.2(v), the sub-pair of the pair $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$ generated by $\mathcal{H}$ is called the hyperpair of $\mathcal{H}.$  [^8] This is the definition in [2], because $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$. It could lead to complications if $\mathcal{P}(\mathcal{H})$ is not distributive, because then the submagma of the power set $\mathcal{P}(\mathcal{H})$ spanned by $\mathcal{H}$ need not be closed under multiplication.
Remark 3.4.
  1. (i)
    Any hyperpair has the important surpassing relation $\subseteq,$ i.e., $S_{1}\preceq S_{2}$ when $S_{1}\subseteq S_{2}.$ Of course $\mathcal{P}(\mathcal{H})$ is a lower semilattice with respect to $(\subseteq)$, but not ${\mathcal{P}}^{\sharp}(\mathcal{H}).$
  2. (ii)
    More generally, if $\mathcal{H}$ has a surpassing relation $\preceq$, then $\mathcal{P}(\mathcal{H})$ has the surpassing relation given by $S_{1}\preceq S_{2}$ if for each $s_{1}\in S_{1}$ there is $s_{2}\in S_{2}$ for which $s_{1}\preceq s_{2}.$
  3. (iii)
    For hypersemigroups $\mathcal{H}$, the condition that $b_{1}\subseteq b_{2}$ implies $f(b_{1})\subseteq f(b_{2})$ is automatic on $\mathcal{P}(\mathcal{H})$.
  4. (iv)
@@
Say a subset $S\subseteq\mathcal{P}(\mathcal{H})$ is weakly neutral if $a\in a\boxplus S$ for each $a\in\mathcal{H}$. The weakly neutral subsets comprise an $\mathcal{H}$-submagma of $\mathcal{P}(\mathcal{H})$.